# Embodied Image Captioning: Self-supervised Learning Agents for Spatially Coherent Image Descriptions

### v1

@@ V-A Exploration policy
We first compare the exploration policies in terms of cosine similarity between the SBERT embeddings of all CoCa captions assigned to each object instance in the corresponding test set (the lower the cosine similarity the better because the descriptions of the same object are semantically different). In Fig. 4 we show the distribution of cosine similarity values in the Gibson test set using CoCa as a captioner. For all policies, 75% of the data shows a cosine similarity lower than 0.6. Random gives the highest similarity at the first, second, and third quartile compared to frontier and CLA. CLA mines 50% of data (median) with similarity values lower than other policies, and the highest density value (number of cosine similarity samples) is almost in correspondence first quartile, while for random and frontier between the first and the second quartile.
@@
Captioner Policy Pseudo-caption $B_{4}$ $M$ $R_{L}$ $CI$ $SP$ $CS$
@@ V-B Pseudo-captioning
We evaluate the performance of pseudo-caption generation using the metrics described in Sec. IV-C and we compare our method with two existing approaches, ECO [12] and IC3 [13] (described in Sec. IV-B) on the corresponding test sets of Gibson and HM3D manually annotated. Results in Table I show that in general our method outperforms both ECO and IC3 across nearly all evaluation metrics in both Gibson and HM3D, for each captioner and policy (values in bold). IC3 has always lower values compared to other methods, due to the fact that the pseudo-caption is a summary of the initial set of captions, hence inaccurate details propagate to the eventual description. Our method has higher $CI$ values, despite ECO uses CIDEr-derived consensus score. In Gibson, the highest improvements of our method compared to ECO can be observed when using CoCa, in the semantic structure of captions ($SP$) with +4.87 percentage points (p.p.) using Random policy and +4 p.p. when using CLA. A similar trend can be observed in HM3D for both CoCa and BLIP2. Table I also shows a comparison of the results based on the policy (values in gray). In particular, mining samples with CLA leads in most cases to higher pseudo-captioning performance for ECO and for our method compared to Random and Frontier policies. Using CLA, the improvement in similarity between predicted and annotated captions ($CS$) for CoCa in Gibson is more than 3 p.p., and for BLIP2 in HM3D the improvement in terms of SPICE ($SP$) is almost 6 p.p.
@@ V-C Captions consistency
Since LD-CPS outperforms the other methods in almost all metrics, we select it as the pseudo-captioning method to fine-tune the captioner. Table II compares the performance of different training procedures and off-the-shelf captioners on the Gibson and HM3D test sets, varying the exploration policy. In general, ChatGPT outperforms all the off-the-shelf captioners. However, the captioning performance of CoCa and BLIP2 improves after fine-tuning, with values also higher than the general purpose model ChatGPT, thanks to the learning of a unique and concise caption generated by our framework. With the majority of combinations of exploration policy and captioner, the use of triplet loss improves the captioning scores compared to standard fine-tuning for CoCa in HM3D and for both CoCa and BLIP2 in Gibson that has image quality lower than HM3D. We show an ablation study of triplet loss weight in Sec. B and qualitative comparisons of the predictions before and after fine-tuning in Sec. C of Supplementary Material.
Captioner Pseudo-caption Policy Fine-tuning $B_{4}$ $M$ $R_{L}$ $CI$ $SP$ $CS$
@@ VI Conclusion
We introduced our modular framework to improve the consistency of captions across different object viewpoints, based on three separate phases: navigation, pseudo-caption generation, and fine-tuning. The experimental validation shows that: the navigation policy explores regions with semantic disagreement among object descriptions higher than baselines; the proposed pseudo-captioning method based on captions frequency and in-context learning, along with the learned policy, generates concise captions having semantic similarity with the human annotation higher than baselines; the triplet loss fine-tuning specializes the caption descriptions to the explored environments while increasing the caption consistency of object instances. Current limitations of the framework lay in generating wrong pseudo-captions in case the initial captions set contains a majority of repeated inaccurate descriptions, and in assuming accurate camera pose and depth estimation during the point cloud generation and segmentation. Future works will focus on extending our approach to dynamic scenarios and on validating the framework on a real robot.
Appendix A Captioners fine-tuning setup
We report the training setup used to fine-tune CoCa [1] and BLIP2 [2] in Table III. Note that we use LoRA [41] to fine-tune BLIP2 due to the large amount of parameters to reduce instability during the fine-tuning processing and reduce the amount of time needed for the fine-tuning processing.
@@ Appendix A Captioners fine-tuning setup
A-A Exploration policy
Figure 8: istribution of cosine similarity between SBERT embeddings of CoCa captions across object instances in HM3D test set, for each exploration policy.
Fig. 8 shows the distribution of captions cosine similarity across the three policies in the HM3D test set. The distributions show a similar trend compared to the ones on the Gibson test set. In this case, frontier gives the highest similarity in all the quartiles. CLA, again, mines 50% of the data with similarity values lower than 0.3. Furthermore, CLA has a higher density in regions with lower similarity in comparison to random and frontier.
Appendix B Ablation study on triplet loss
Table IV shows the effect of the $\lambda_{tr}$ coefficient on the captioners fine-tuning. We experiment $\lambda_{tr}=\{1,0.5,0.1\}$. Results show that a choice of $\lambda_{tr}=0.1$ achieves higher performance values compared to standard fien-tuning and other triplet weights. With $\lambda_{tr}=1$ the fine-tuned model has worse performance values compared to other triplet loss weights and standard fine-tuned model with all the exploration policies apart from CLA.
Captioner Pseudo-caption Policy Fine-tuning $B_{4}$ $M$ $R_{L}$ $CI$ $SP$ $CS$
@@ Appendix B Ablation study on triplet loss
TABLE IV: Comparison of captioning performance on Gibson and HM3D testing sets, varying fine-tuning method and policy.
@@
Appendix C Ablation study on using ground truth object detections
We investigate the effect of the errors introduced using Mask2Former instance segmentation model in our framework. We train the whole pipeline on the subset of the Gibson dataset containing semantic object labels. In particular, we replace the output of the method in Section III-A with the ground truth object IDs. Note that this set is smaller than the one used in the rest of the paper, consisting of 25 train scenes and 5 validation scenes. Table V shows the results of this ablation study. Fine-tuning with semantic annotations leads to overall performance improvements compared to training without fine-tuning, while triplet loss provides additional gains. The trends remain consistent with those observed using the real object detector, suggesting that the pipeline remains effective even in the presence of segmentation noise.
Captioner Pseudo-caption Policy Fine-tuning $B_{4}$ $M$ $R_{L}$ $CI$ $SP$ $CS$
Gibson Semantic CoCa - - NoFT 12.15 21.47 42.83 0.63 28.30 63.30
Ours Random Standard 14.48 23.40 51.38 0.93 33.37 72.10
Triplet 14.75 24.20 51.29 0.94 36.50 71.41
Frontier Standard 13.11 21.56 46.67 0.80 31.33 67.63
Triplet 13.48 21.82 45.82 0.76 29.38 66.28
CLA Standard 13.29 21.37 47.87 0.74 31.39 64.92
Triplet 13.56 22.01 48.23 0.81 34.14 66.09
TABLE V: Comparison of captioning performance on Gibson testing sets of model trained using Gibson semantic annotations, varying fine-tuning method and policy.
@@ Appendix D Qualitative comparison of captioners prediction
In Fig. 7 we present some qualitative examples of predicted captions before and after fine-tuning (with triplet loss) for both CoCa and BLIP2. Off-the-shelf models predict captions containing wrong attributes (colors) or object categories not present in the images e.g. cat, bird’s feathers, television. On the contrary, the fine-tuned models show higher accuracy in describing details of objects and generate predictions that are more consistent across different views of the same object, showing that learning the unique pseudo-caption for each object instance combined with the triplet loss enhances captions accuracy and consistency.
@@ Appendix E Pseudo-captioning and ChatGPT prompts
Listing 1 shows the proposed text prompt for the pseudo-caption generation, where str(captions_freq_list) is a list of captions for different views of the same object with their frequency. Our approach combines the frequency of the captions with in-context learning (provided examples of how to solve the task), to account for wrong or inconsistent captions.
Listing 2 shows the text prompt used to generate results with ChatGPT [4]. We use ChatGPT to generate captions from the objects crop.
Listing 1: The prompt for our proposed method.
@@
Listing 2: The prompt for off-the-shelf ChatGPT.
Please provide an image caption for the provided picture. Do not
use more than 5-7 words to describe the object in the image.
Use simple words and dont use more than one adjective per
noun. Some examples: A red couch with pillows on it’,’A
television set on top of a table’,’A grey patterned armchair

### v2

@@ V-A Exploration policy
We first compare the exploration policies in terms of cosine similarity between the SBERT embeddings of all CoCa captions assigned to each object instance in the corresponding test set (the lower the cosine similarity the better because the descriptions of the same object are semantically different). In Fig. 3 we show the distribution of cosine similarity values in the Gibson test set using CoCa as a captioner. For all policies, 75% of the data shows a cosine similarity lower than 0.6. Random gives the highest similarity at the first, second, and third quartile compared to frontier and CLA. CLA mines 50% of data (median) with similarity values lower than other policies, and the highest density value (number of cosine similarity samples) is almost in correspondence first quartile, while for random and frontier between the first and the second quartile. Note that CLA shows the highest value of cosine similarity due to the fact that the policy can initially visit also objects with consistent captions, in order to build the disagreement map.
@@
Captioner Policy Pseudo-caption $B_{4}$ $M$ $R_{L}$ CICI SPSP CSCS
@@ V-B Pseudo-captioning
We evaluate the performance of pseudo-caption generation using the metrics described in Sec. IV-C and we compare our method with two existing approaches, ECO [12] and IC3 [13] (described in Sec. IV-B) on the corresponding test sets of Gibson and HM3D manually annotated. Results in Table I show that in general our method outperforms both ECO and IC3 across nearly all evaluation metrics in both Gibson and HM3D, for each captioner and policy (values in bold). IC3 has always lower values compared to other methods, due to the fact that the pseudo-caption is a summary of the initial set of captions, hence inaccurate details propagate to the eventual description. Moreover, ECO depends on CLIP models that might select the wrong pseudo-caption to associate with the image. Our method shows higher $CI$ values, despite ECO uses CIDEr-derived consensus score. In Gibson, the highest improvements of our method compared to ECO can be observed when using CoCa, in the semantic structure of captions ($SP$) with +4.87 percentage points (p.p.) using Random policy and +4 p.p. when using CLA. A similar trend can be observed in HM3D for both CoCa and BLIP2. Table I also shows a comparison of the results based on the policy (values in gray). In particular, mining samples with CLA leads in most cases to higher pseudo-captioning performance for ECO and for our method compared to Random and Frontier policies, thanks to the maximization of the disagreement map. Using CLA, the improvement in similarity between predicted and annotated captions ($CS$) for CoCa in Gibson is more than 3 p.p., and for BLIP2 in HM3D the improvement in terms of SPICE ($SP$) is almost 6 p.p.
@@ V-C Captions consistency
Since LD-CPS outperforms the other methods in almost all metrics, we select it as the pseudo-captioning method to fine-tune the captioner. Table II compares the performance of different training procedures and off-the-shelf captioners on the Gibson and HM3D test sets, varying the exploration policy. In general, ChatGPT outperforms all the off-the-shelf captioners, yet the captioning performance of CoCa and BLIP2 improves after fine-tuning, with values also higher than the general purpose model ChatGPT, thanks to the learning of a unique and concise caption generated by our framework. With the majority of combinations of exploration policy and captioner, the use of triplet loss, which forces predictions of different object views to be similar, improves the captioning scores compared to standard fine-tuning in both Gibson and HM3D. We show an ablation study of triplet loss weight in Sec. B and qualitative comparisons of the predictions before and after fine-tuning in Sec. D of Supp. Mat.
Captioner Pseudo-caption Policy Fine-tuning $B_{4}$ $M$ $R_{L}$ CICI SPSP CSCS
@@ VI Conclusion
We introduced our modular framework to improve the consistency of captions across different object viewpoints, based on three separate phases: navigation, pseudo-caption generation, and fine-tuning. The experimental validation shows that: the navigation policy explores regions with semantic disagreement among object descriptions higher than baselines; the proposed pseudo-captioning method based on captions frequency and in-context learning, along with the learned policy, generates concise captions having semantic similarity with the human annotation higher than baselines; the triplet loss fine-tuning specializes the caption descriptions to the explored environments while increasing the caption consistency of object instances.
Limitations. Current limitations of the framework lay in generating wrong pseudo-captions in case the initial captions set contains a majority of repeated inaccurate descriptions, due to the proposed prompt based on caption frequency, and in aggregating inaccurate object instances in the point cloud due to the assumption of accurate camera pose and depth estimation during the point cloud generation and segmentation. Moreover, the framework is not optimized to run real-time on a standalone robot. Future works will focus on extending our approach to dynamic scenarios and on validating the framework on a real robot.
Acknowledgement
We acknowledge the financial support from PNRR MUR Project PE0000013 ”Future Artificial Intelligence Research (hereafter FAIR)”, funded by the European Union – NextGenerationEU, CUP J53C22003010006. This project has received funding from the European Union’s Horizon research and innovation programme G.A. n. 101070227 (CONVINCE).
Appendix A Experimental setup
Datasets. Gibson reflects the complexity of 572 real-world environments featuring object categories such as tables, sofas, plants, in different settings (rooms and object instances). HM3D comprises 1,000 reconstructed indoor residential and commercial environments having higher quality than Gibson [32].
Methods under comparison. We compare three exploration policies: random goals, frontier exploration and the learned policy (CLA). The first two policies use the same path planner that first converts the explored map into a visibility graph (nodes represent feasible goals and edges the straight trajectory without obstacles) via skeletonization, then computes the optimal trajectory between the current position and the goal using graph search, considering nodes in the trajectory as sub-goals. Moreover, we delete repeated samples caused by the agent revisiting certain poses during exploration, to train and test on different images. Training images are augmented with 50% probability using horizontal flip, gaussian noise and affine transformations.
Captioners fine-tuning setup. We report the training setup used to fine-tune CoCa [1] and BLIP2 [2] in Table III. Note that we use LoRA [45] to fine-tune BLIP2 due to the large amount of parameters to reduce instability during the fine-tuning processing and reduce the amount of time needed for the fine-tuning processing.
@@ Appendix A Experimental setup
Appendix B Ablation studies
Pseudo-caption ablation. Results in Tab. IV show that compared to LLaMa, Mistral achieves -1.85 p.p. in SP and -0.42 p.p. in CI while Qwen achieves -1.41 p.p. in SP and -0.13 p.p. in CI using BLIP2. Also using CoCa, LLaMa outperforms other LLMs. Varying the LLMs, our prompt with caption frequency outperforms ECO and IC3 (Tab.1 main paper).
Captioner LLM $B_{4}$ $M$ $R_{L}$ CICI SPSP CSCS
CoCa Mistral-7B 17.21 22.07 44.70 1.13 29.29 71.63
Qwen3-8B 19.95 23.95 49.60 1.31 33.59 73.37
BLIP-2 Mistral-7B 16.46 24.05 48.02 1.20 33.73 72.11
Qwen3-8B 19.37 24.62 51.43 1.49 34.17 73.02
TABLE IV: LD-CPS ablation on HM3D with CLA.
Ablation study on triplet loss. Table V shows the effect of the $\lambda_{tr}$ coefficient on the captioners fine-tuning. We experiment $\lambda_{tr}=\{1,0.5,0.1\}$. Results show that a choice of $\lambda_{tr}=0.1$ achieves higher performance values compared to standard fien-tuning and other triplet weights. With $\lambda_{tr}=1$ the fine-tuned model has worse performance values compared to other triplet loss weights and standard fine-tuned model with all the exploration policies apart from CLA.
Captioner Pseudo-caption Policy Fine-tuning $B_{4}$ $M$ $R_{L}$ CICI SPSP CSCS
@@ Appendix B Ablation studies
TABLE V: Comparison of captioning performance on Gibson and HM3D testing sets, varying fine-tuning method and policy.
@@
Appendix C Inference time
On an NVIDIA Tesla V100 16GB, the time for an exploration step is 0.266 s for CLA equipped with CoCa, 0.271 s equipped with BLIP2 (averaged over 300 steps); of these timings, the captioning inference time is 6.65 ms for CoCa and 10.2 ms for BLIP2 (averaged over 754 bounding boxes); and the pseudo-captioning time per object instance is 1.02 s (averaged over all 160 testing objects).
@@ Appendix D Qualitative comparison of captioners prediction
In Fig. 6 we present some qualitative examples of predicted captions before and after fine-tuning (with triplet loss) for both CoCa and BLIP2. Off-the-shelf models predict captions containing wrong attributes (colors) or object categories not present in the images, e.g., cat, bird’s feathers, television. On the contrary, the fine-tuned models show higher accuracy in describing details of objects and generate predictions that are more consistent across different views of the same object, showing that learning the unique pseudo-caption for each object instance combined with the triplet loss enhances captions accuracy and consistency.
@@ Appendix E Pseudo-captioning and ChatGPT prompts
Listing 1 shows the text prompt used to generate results with ChatGPT [4]. We use ChatGPT to generate captions from the objects crop.
Listing 2 shows the proposed text prompt for the pseudo-caption generation, where str(captions_freq_list) is a list of captions for different views of the same object with their frequency. Our approach combines the frequency of the captions with in-context learning (provided examples of how to solve the task), to account for wrong or inconsistent captions.
Listing 1: The prompt for off-the-shelf ChatGPT.
Please provide an image caption for the provided picture. Do not
use more than 5-7 words to describe the object in the image. Use
simple words and dont use more than one adjective per noun.
Some examples: A red couch with pillows on it’,’A television
set on top of a table’,’A grey patterned armchair
Listing 2: The prompt for our proposed method.
@@
Appendix F Performance on occluded objects
Tab. VI reports a comparison between off-the-shelf and fine-tuned captioners (using triplet loss) on a manually collected dataset of 50 images (25 from Gibson and 25 from HM3D) of objects, with 5 images per dataset for each category in the main paper, under varying levels of occlusion based on visual inspection. The performance is evaluated using standard captioning metrics (BLEU-4, METEOR, ROUGE-L, CIDEr, SPICE and cosine similarity), and results are averaged across Gibson and HM3D. The results show that our fine-tuned models outperform off-the-shelf captioners for both CoCa and BLIP-2, indicating improved robustness to partial views and occlusions. Qualitative examples in Fig. 7 illustrate the increased consistency and accuracy of the generated captions in challenging visual conditions.
Env Finetuning $B_{4}$ $M$ $R_{L}$ CICI SPSP CSCS
CoCa NoFT 16.01 22.19 46.37 0.88 31.07 68.71
Triplet 17.95 24.10 48.20 1.05 32.98 74.03
BLIP-2 NoFT 20.85 23.18 50.39 1.32 36.48 73.92
Triplet 22.70 25.29 53.12 1.38 39.90 77.58
TABLE VI: Comparison of captioning performance on occluded objects. KEYS – $B_{4}$: BLEU, $M$: METEOR, $R_{L}$: ROUGE-L, $CI$: CIDER, $SP$: SPICE, $CS$: cosine similarity between SBERT embedding of predicted and annotated captions.
Figure 7: Captioning with different levels of occlusion. C: CoCa; B: BLIP2; F: fine-tuned (triplet loss)